OpenStudy (anonymous):

Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation.

OpenStudy (anonymous):

@bailey_ann2398 : try this questn from a easy cotext it is give Meg has a Can containing 80 % orange juice so 20 % will be water let the quantity of juice (mixed with water) be x ltr.. so 20 % of x is 1 ltr ie 20/100 \[(\frac{ 20 }{ 100 } \times X = 1 ltr) \]

OpenStudy (anonymous):

so x/5=1ltr ie x= 5 ltr total juice is 5 ltr ie (4 ltr orange juice and 1 ltr water)

OpenStudy (anonymous):

@bailey_ann2398 : hope you got it and can try the rest...

OpenStudy (anonymous):

thanks for the help :) @bhaskar.pathk

OpenStudy (anonymous):

@bailey_ann2398 : :) Always Wlcm